# DrivAerNet++: A Large-Scale Multimodal Car Dataset with Computational Fluid Dynamics Simulations and Deep Learning Benchmarks

### v1

@@ Abstract
We present DrivAerNet++, the largest and most comprehensive multimodal dataset for aerodynamic car design. DrivAerNet++ comprises 8,000 diverse car designs modeled with high-fidelity computational fluid dynamics (CFD) simulations. The dataset includes diverse car configurations such as fastback, notchback, and estateback, with different underbody and wheel designs to represent both internal combustion engines and electric vehicles. Each entry in the dataset features detailed 3D meshes, parametric models, aerodynamic coefficients, and extensive flow and surface field data, along with segmented parts for car classification and point cloud data. This dataset supports a wide array of machine learning applications including data-driven design optimization, generative modeling, surrogate model training, CFD simulation acceleration, and geometric classification. With more than 39 TB of publicly available engineering data, DrivAerNet++ fills a significant gap in available resources, providing high-quality, diverse data to enhance model training, promote generalization, and accelerate automotive design processes. Along with rigorous dataset validation, we also provide ML benchmarking results on the task of aerodynamic drag prediction, showcasing the breadth of applications supported by our dataset. This dataset is set to significantly impact automotive design and broader engineering disciplines by fostering innovation and improving the fidelity of aerodynamic evaluations.
@@ 2 Related work
Usama et. al 2021 [68] 500 40 (2D) P
Li et. al 2023  [43]* 551 6 (3D) M,P,C
Rios et. al 2021 [58$\dagger$ 600 - M,PC
Li et. al 2023 [43$\dagger$ 611 - M,C
Umetani et. al 2018 [67$\dagger$ 889 - M,C
Gunpinar et. al 2019 [29] 1,000 21 (2D) P
Jacob et. al 2021 [37]* 1,000 15 (3D) M,C,P
Trinh et. al 2024 [66] 1,121 - M,C
Remelli et. al 2020 [55$\dagger$ 1,400 - M,C
Baque el al. 2018 [6] 2,000 21 (3D) M,P
Song et. al 2023 [61$\dagger$ 2,474 - M
@@ 6 Conclusion
In this paper, we introduced DrivAerNet++, the largest, multimodal 3D dataset for data-driven aerodynamic design, incorporating high-fidelity CFD simulations and a diverse array of car designs. Our dataset includes 8,000 cars based on industry-standard shapes, offering extensive coverage of various aerodynamic performance metrics. The dataset, requiring 39 TB of storage, is significantly larger than comparable datasets in engineering and is made publicly available. Additionally, the computational cost of generating DrivAerNet++ was an order of magnitude larger than that of the recently published CFD dataset from [43], which utilized 185,744 CPU-hours, whereas ours required 3,000,000 CPU-hours.
This comprehensive dataset supports a wide range of machine learning tasks, including surrogate modeling of aerodynamic performance, acceleration of CFD simulations, data-driven design optimization, generative AI, shape and part classification, and 3D shape reconstruction. We also presented the first benchmarking results, demonstrating the effectiveness of geometric deep learning models and AutoML frameworks for drag coefficient prediction. Additionally, we explored the challenges of creating a generalized model for surrogate modeling of drag across different car categories. While models trained on a single car category performed well, their performance in terms of $R^{2}$ decreased significantly from 0.82-0.9 to 0.6 when applied to the full diverse dataset. This underscores the complexity of achieving robust performance across varied designs.
@@ 7 Limitations and future work
The DrivAerNet++ dataset, with its 8,000 car designs, covers most conventional car designs. However, the simulation fidelity is slightly lower than what is typically used in the industry (CFD meshes with $\mathcal{O}$(100M) cells [4]). Additionally, the highly turbulent and time-dependent flow around a 3D car introduces errors when using steady-state RANS simulations. While the $k$-$\omega$-SST model [45] generally provides accurate results, it struggles with predicting flow separation and reattachment [36, 4]. Future work should involve hybrid RANS-LES methods to better capture the time-dependent nature of the flow field.
Moreover, the surrogate models we trained are not complex enough to learn the intricate geometric and aerodynamic features. More advanced models, such as Geometry-Informed Neural Operators [43], Convolutional Occupancy Networks [51], and sophisticated graph models, should be tested. We primarily focused on the surrogate modeling of drag since it is the most critical factor in the initial design phase. However, leveraging DrivAerNet++ for other tasks, such as accelerating CFD simulations, is necessary for a more comprehensive approach.
To enhance the dataset, future work will focus on integrating transient CFD simulations, and incorporating additional modalities like 2D image renderings and multimodal learning approaches. This will improve models accuracy and robustness, driving innovation in automotive design and optimization.
@@
The authors thank Ravi Nimbalkar and Yatin Kumbhar from BETA CAE Systems USA Inc for their support in creating the parametric models in ANSA®. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC, and consultation resources that have contributed to the research results reported within this paper. Special thanks to Lauren Milechin for her support in setting up the simulation environment on MIT SuperCloud.
@@
Appendix A Broader Impact
@@
Appendix B DrivAerNet++ Dataset Generation
B.1 Generating a wide range of car designs
@@ B.2 CFD Meshing
B.3 Numerical Simulation
B.3.1 Domain and Boundary Conditions
For the CFD simulations, we used the open-source software OpenFOAM® v11.444https://openfoam.org/version/11/, which is a comprehensive suite of C++ modules that allows for the customization of solvers and utilities in CFD studies and supports parallelization. We utilized 1:1 scaled models of all car designs and the SIMPLE algorithm (Semi-Implicit Method for Pressure Linked Equations) to couple pressure and velocity in the incompressible fluid simulations. These were carried out using the simpleFoam solver, which is ideal for steady-state, turbulent, and incompressible flow simulations (In OpenFOAM® v11, the incompressibleFluid modular solver replicates simpleFoam).
@@ B.3.1 Domain and Boundary Conditions
We used the potentialFoam solver to initialize the OpenFOAM® simulations. This solver calculates a potential flow solution to quickly establish an initial velocity field, thereby accelerating the convergence of the solution in complex aerodynamic simulations [28]. The potential flow assumption neglects viscous effects and assumes irrotational flow, providing a good starting point for the simpleFoam solver.
@@
The simulations were conducted using air with the properties listed in Table 5.
B.3.2 Validation of the Numerical Results
In this section, we present validation results for the baseline simulations. The assumption is that given a baseline geometry, conducting the CFD simulation and benchmarking the results against both reference simulations and wind tunnel experimental data should establish a foundation for using the same CFD meshing approach, boundary conditions, and solvers for the morphed geometries of this baseline geometry. However, DrivAerNet++ encompasses various car designs and categories, making a single mesh unsuitable due to variations in car length, rear end, and underbody features. Therefore, as detailed in the meshing section B.2, we employ different meshing strategies based on car categories. Similarly, we benchmark the CFD simulation results across different geometries, focusing on the following car categories:
@@ B.4 Convergence Detection
B.5 Scaling Up the Dataset Size vs Simulation Fidelity
Appendix C In-Depth Insights from DrivAerNet++
@@ C.1 Dataset Annotations
Appendix D Deep Surrogate Models
D.1 RegDGCNN
@@ D.2 PointNet
We also considered another variant of PointNet for the task of pressure prediction on the car surface (see Section F.1). The PointNet model for pressure prediction consists of an input layer for 3D coordinates, followed by three convolutional layers with increasing channel sizes: 64, 128, and 1024. A Spatial Transformer Network (STN) [38] is used to align the input points, enhancing the network’s ability to learn spatial features.
D.3 Graph Neural Network
@@ Appendix E Evaluation
Appendix F Additional Experiments
@@
F.1 3D Surface Field Prediction
@@
F.2 t-SNE Visualization of the Design Space
In this section, we train a PointNet [53] model for 3D shape classification and utilize the learned features to visualize the design space. By reducing the dimensionality of these features to 2D using t-SNE (t-distributed stochastic neighbor embedding) [69], we can create a plot that illustrates the similarity and variation between different car designs. This visualization aids designers in linking new designs to existing ones, potentially identifying the most similar designs and their performance using methods like K-Nearest Neighbors (KNN) [52].
Additionally, using t-SNE to visualize the high-dimensional design space provides an efficient means to explore the vast design space. This method eliminates the need to compute complex metrics, such as the Chamfer distance, between a new design and all existing designs. By performing inference on a new design with a pretrained model, extracting the learned features, reducing their dimensionality, and visualizing their position in 2D space, we can easily assess how similar or different a new design is compared to existing ones. Figure 20 demonstrates the t-SNE visualization, showing distinguished clusters for different car categories: Notchback, Estateback, and Fastback. Each point in the plot represents a unique design, illustrating the ability of t-SNE to capture and separate the underlying differences in car shapes effectively. In general, t-SNE is a powerful tool for reducing high-dimensional data to two or three dimensions while preserving the relative distances between data points, which helps in identifying patterns and similarities.
Additionally, incorporating information about the aerodynamic performance of specific designs is crucial for analyzing new designs. Although Figure 20 focuses on shape similarities, the methodology can be extended to include performance values, providing a more comprehensive understanding of different designs. This combined approach enables a more thorough analysis of design efficiency and performance, facilitating better-informed decisions in the design process.

### v2

@@ Abstract
We present DrivAerNet++, the largest and most comprehensive multimodal dataset for aerodynamic car design. DrivAerNet++ comprises 8,000 diverse car designs modeled with high-fidelity computational fluid dynamics (CFD) simulations. The dataset includes diverse car configurations such as fastback, notchback, and estateback, with different underbody and wheel designs to represent both internal combustion engines and electric vehicles. Each entry in the dataset features detailed 3D meshes, parametric models, aerodynamic coefficients, and extensive flow and surface field data, along with segmented parts for car classification and point cloud data. This dataset supports a wide array of machine learning applications including data-driven design optimization, generative modeling, surrogate model training, CFD simulation acceleration, and geometric classification. With more than 39 TB of publicly available engineering data, DrivAerNet++ fills a significant gap in available resources, providing high-quality, diverse data to enhance model training, promote generalization, and accelerate automotive design processes. Along with rigorous dataset validation, we also provide ML benchmarking results on the task of aerodynamic drag prediction, showcasing the breadth of applications supported by our dataset. This dataset is set to significantly impact automotive design and broader engineering disciplines by fostering innovation and improving the fidelity of aerodynamic evaluations. Dataset and code available at: https://github.com/Mohamedelrefaie/DrivAerNet
@@ 2 Related Work
Usama et al. 2021 [70] 500 40 (2D) P
Li et al. 2023  [45]* 551 6 (3D) M,P,C
Rios et al. 2021 [60$\dagger$ 600 - M,PC
Li et al. 2023 [45$\dagger$ 611 - M,C
Umetani et al. 2018 [69$\dagger$ 889 - M,C
Gunpinar et al. 2019 [31] 1,000 21 (2D) P
Jacob et al. 2021 [39]* 1,000 15 (3D) M,C,P
Trinh et al. 2024 [68] 1,121 - M,C
Remelli et al. 2020 [57$\dagger$ 1,400 - M,C
Baque et al. 2018 [7] 2,000 21 (3D) M,P
Song et al. 2023 [63$\dagger$ 2,474 - M
@@ 3 Dataset Presentation
Comparison of mesh resolutions across large-scale datasets
Here we compare the mesh resolution on car surfaces across different existing large-scale datasets. The results are summarized in Table 2. The mesh resolution comparison highlights the varying levels of detail used in different large-scale datasets for aerodynamic simulations. DrivAerNet++ (ours) provides the highest mesh resolution, ranging from 350,000 to 750,000 cells, ensuring detailed representation of car surfaces for accurate simulations. In contrast, other datasets like the Ahmed body by Li et al. (2023) [45] and ShapeNet (Song et al., 2023 [63]) use lower resolutions, with 100,000 cells and 30,000–50,000 cells, respectively. This finer mesh resolution in DrivAerNet++ enables better fidelity in aerodynamic analyses, particularly for capturing intricate details crucial for engineering design.
Table 2: Comparison of mesh resolutions on car surfaces in large-scale datasets.
Dataset Mesh Resolution on Car Surface
DrivAerNet++ (Ours) 350k to 750k cells
Ahmed body (Li et al., 2023) [45] 100k cells
ShapeNet (Song et al., 2023) [63] 30–50k cells
ShapeNet (Umetani et al., 2018) [69] 10k cells
@@ 6 Conclusion
In this paper, we introduced DrivAerNet++, the largest, multimodal 3D dataset for data-driven aerodynamic design, incorporating high-fidelity CFD simulations and a diverse array of car designs. Our dataset includes 8,000 cars based on industry-standard shapes, offering extensive coverage of various aerodynamic performance metrics. The dataset, requiring 39 TB of storage, is significantly larger than comparable datasets in engineering and is made publicly available. The computational cost of generating DrivAerNet++ was an order of magnitude larger than that of the recently published CFD dataset from [45], which utilized 185,744 CPU-hours, whereas ours required 3,000,000 CPU-hours.
This comprehensive dataset supports a wide range of machine learning tasks, including surrogate modeling of aerodynamic performance, acceleration of CFD simulations, data-driven design optimization, generative AI, shape and part classification, and 3D shape reconstruction. We also presented the first benchmarking results, demonstrating the effectiveness of geometric deep learning models and AutoML frameworks for drag coefficient prediction. Furthermore, we explored the challenges of creating a generalized model for surrogate modeling of drag across different car categories. While models trained on a single car category performed well, their performance in terms of $R^{2}$ decreased significantly from 0.82-0.9 to 0.6 when applied to the full diverse dataset. This underscores the complexity of achieving robust performance across varied designs.
@@ 7 Limitations and Future Work
The surrogate models we trained on DrivAerNet++ are currently not complex enough to fully capture the intricate geometric and aerodynamic features present in diverse car designs. More advanced models, such as Geometry-Informed Neural Operators [45], Convolutional Occupancy Networks [53], and sophisticated graph models, should be explored to improve predictive performance. We primarily focused on the surrogate modeling of drag, as it is a critical factor in the initial design phase. However, leveraging DrivAerNet++ for additional applications, such as accelerating CFD simulations, would provide a more comprehensive tool for aerodynamic design.
To enhance the dataset, future work will focus on integrating transient CFD simulations and incorporating additional modalities, such as 2D image renderings and multimodal learning approaches. These enhancements aim to improve model accuracy and robustness, fostering further innovation in automotive design and optimization.
@@
The authors thank Ravi Nimbalkar and Yatin Kumbhar from BETA CAE Systems USA Inc for their support in creating the parametric models in ANSA®. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC and consultation resources that have contributed to the research results reported within this paper. Special thanks to Chris Hill for providing access to MIT SuperCloud and to Lauren Milechin for her support in setting up the simulation environment. Additionally, the authors wish to thank Amy Nurnberger from MIT and Sonia Barbosa from Harvard University for their support with uploading the dataset to Harvard Dataverse. Finally, the authors extend their gratitude to Leonid Andreev for his technical support with uploading the dataset to Harvard Dataverse.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] This is done in Section 7
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] Our work does not involve human participants or crowd-sourcing. The data used is licensed, consented, and adheres to the FAIR (Findable, Accessible, Interoperable, and Reusable) principles [73].
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] The source code utilized for model training and assessment in this research can be accessed at the given repository https://github.com/Mohamedelrefaie/DrivAerNet, where the README file offers detailed guidance for usage.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] DrivAerNet++ is designed as a benchmark dataset for 3D problems, complete with a split into training, validation, and test sets to ensure fair comparisons among various machine and deep learning models. Details on the dataset splits can be found here: https://github.com/Mohamedelrefaie/DrivAerNet
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] The high-fidelity CFD simulations for DrivAerNet++ were conducted on the MIT SuperCloud, utilizing parallelization across 60 nodes, totaling 2880 CPU cores. The simulations took approximately 3 × 106 CPU-hours to complete. For more details, refer to Section 3: Computational cost.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] The baseline CAD geometry of the DrivAer model is available online for free download [49], but registration is required, and we have cited the creators.
    2. (b)
      Did you mention the license of the assets? [N/A] The CAD geometry is not provided under a license, however it requires registration for access.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] All new car designs created by the authors of this paper are included in the dataset and will be hosted on Harvard Dataverse Repository.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] Consent is not applicable as all car designs were created by the authors of this paper.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A] The data does not contain any personally identifiable information or offensive content.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@
Appendix A Reproducibility Statement
We provide a GitHub repository to facilitate the reproduction of all experiments in the main paper, along with links to download the dataset. The repository includes the following: code to replicate the machine learning experiments and code to generate the figures presented in the paper. The full dataset will be made available after the peer-review process is complete. The training and experiments were performed on a machine equipped with AMD EPYC 7763 64-Core Processors, totaling 256 CPU cores, and four Nvidia A100 80GB GPUs.
Appendix B URL and Links
DrivAerNet++ Dataset: The DrivAerNet [24] and the DrivAerNet++ datasets are available on our GitHub repository at https://github.com/Mohamedelrefaie/DrivAerNet.
Croissant Metadata: The URL to Croissant metadata record documenting the dataset/benchmark is available on our GitHub repository at https://github.com/Mohamedelrefaie/DrivAerNet/tree/main/mlcroissant.
DL Code: The source code utilized for model training and assessment in this research can be accessed at the given repository https://github.com/Mohamedelrefaie/DrivAerNet, where the README file offers detailed guidance for usage.
Tutorials and Browsing Tools: Additional resources, including data handling tutorials and contribution guidelines, are available at https://github.com/Mohamedelrefaie/DrivAerNet/tree/main/tutorials
Appendix C Author Statement and Data License Confirmation
We, the authors, hereby declare that we bear full responsibility for any violations of rights or other issues arising from the use of the data. We confirm that the dataset is available under the Creative Commons Attribution-NonCommercial (CC BY-NC) license. This license allows others to remix, tweak, and build upon our work non-commercially, and although their new works must also acknowledge our original work and be non-commercial, they don’t have to license their derivative works on the same terms.
Appendix D Broader Impact
@@
Appendix E DrivAerNet++ Dataset Generation
E.1 Generating a wide range of car designs
@@ E.2 CFD Meshing
E.3 Numerical Simulation
E.3.1 Domain and Boundary Conditions
For the CFD simulations, we used the open-source software OpenFOAM® v11, [^4] https://openfoam.org/version/11/ a comprehensive suite of C++ modules that supports solver customization, utility development, and parallelization in CFD studies. In the field of aerodynamics, only two open-source tools—OpenFOAM [30] and SU2 [21]—are commonly used, with OpenFOAM being the state-of-the-art tool most widely adopted in both industry and academia for car aerodynamics. A review of the literature indicates no significant technical advantages of using SU2 over OpenFOAM for this application. We utilized 1:1 scaled models of all car designs and the SIMPLE algorithm (Semi-Implicit Method for Pressure Linked Equations) to couple pressure and velocity in the incompressible fluid simulations. These were carried out using the simpleFoam solver, which is ideal for steady-state, turbulent, and incompressible flow simulations (In OpenFOAM® v11, the incompressibleFluid modular solver replicates simpleFoam).
@@ E.3.1 Domain and Boundary Conditions
We used the potentialFoam solver to initialize the OpenFOAM® simulations. This solver calculates a potential flow solution to quickly establish an initial velocity field, thereby accelerating the convergence of the solution in complex aerodynamic simulations [30]. The potential flow assumption neglects viscous effects and assumes irrotational flow, providing a good starting point for the simpleFoam solver. The simulations were conducted using air with the properties listed in Table 6.
E.3.2 Meshing Sensitivity
Mesh sensitivity analysis is conducted to ensure that simulation results are independent of the mesh resolution, which can significantly impact the accuracy of the CFD simulations. By varying the mesh density and observing its effect on outputs like the drag coefficient, analysts can determine the optimal mesh size that balances computational cost with solution accuracy. This process helps to confirm that the results are not artifacts of the chosen mesh and improves confidence in the reliability of the simulation predictions. We conducted a mesh sensitivity analysis to examine the effect of mesh resolution on the drag coefficient specifically for the fastback configuration. The findings are summarized in Table LABEL:tab:mesh_resolution, which shows that as the mesh quality improves from coarse (6M cells) to fine (24M cells), the relative error in drag coefficient decreases from 8.21% to 2.17%, indicating a convergence towards the reference CFD results from TUM [34]. However, the computational time increases significantly with finer meshes, demonstrating the trade-off between accuracy and computational cost. This analysis highlights the importance of selecting an appropriate mesh resolution to balance precision with efficiency in CFD simulations.
Table 7: Influence of mesh resolution on drag coefficient for the fastback case. The relative error is computed with respect to the CFD results from TUM [34].
Mesh Quality Number of Cells Relative Time Relative Error in Drag (%)
Coarse 6M 1.00 8.21%
Medium 12M 1.25 6.55%
Fine 24M 1.88 2.17%
E.3.3 Validation of the Numerical Results
In this section, we present validation results for the baseline simulations. The assumption is that given a baseline geometry, conducting the CFD simulation and benchmarking the results against both reference simulations and wind tunnel experimental data should establish a foundation for using the same CFD meshing approach, boundary conditions, and solvers for the morphed geometries of this baseline geometry. However, DrivAerNet++ encompasses various car designs and categories, making a single mesh unsuitable due to variations in car length, rear end, and underbody features. Therefore, as detailed in the meshing section E.2, we employ different meshing strategies based on car categories. Similarly, we benchmark the CFD simulation results across different geometries, focusing on the following car categories:
@@ E.4 Convergence Detection
Appendix F In-Depth Insights from DrivAerNet++
@@ F.1 Dataset Annotations
Appendix G Deep Surrogate Models
G.1 RegDGCNN
@@ G.2 PointNet
We also considered another variant of PointNet for the task of pressure prediction on the car surface (see Section I.1). The PointNet model for pressure prediction consists of an input layer for 3D coordinates, followed by three convolutional layers with increasing channel sizes: 64, 128, and 1024. A Spatial Transformer Network (STN) [40] is used to align the input points, enhancing the network’s ability to learn spatial features.
G.3 Graph Neural Network
@@ Appendix H Evaluation
Appendix I Additional Experiments
@@
I.1 3D Surface Field Prediction
@@
I.2 t-SNE Visualization of the Design Space
In this section, we train a PointNet [55] model for 3D shape classification and utilize the learned features to visualize the design space. By reducing the dimensionality of these features to 2D using t-SNE (t-distributed stochastic neighbor embedding) [71], we can create a plot that illustrates the similarity and variation between different car designs. This visualization aids designers in linking new designs to existing ones, potentially identifying the most similar designs and their performance using methods like K-Nearest Neighbors (KNN) [54]. Using t-SNE to visualize the high-dimensional design space provides an efficient means to explore the vast design space. This method eliminates the need to compute complex metrics, such as the Chamfer distance, between a new design and all existing designs. By performing inference on a new design with a pretrained model, extracting the learned features, reducing their dimensionality, and visualizing their position in 2D space, we can easily assess how similar or different a new design is compared to existing ones. Figure 20 demonstrates the t-SNE visualization, showing distinguished clusters for different car categories: Notchback, Estateback, and Fastback. Each point in the plot represents a unique design, illustrating the ability of t-SNE to capture and separate the underlying differences in car shapes effectively. In general, t-SNE is a powerful tool for reducing high-dimensional data to two or three dimensions while preserving the relative distances between data points, which helps in identifying patterns and similarities.
Furthermore, incorporating information about the aerodynamic performance of specific designs is crucial for analyzing new designs. Although Figure 20 focuses on shape similarities, the methodology can be extended to include performance values, providing a more comprehensive understanding of different designs. This combined approach enables a more thorough analysis of design efficiency and performance, facilitating better-informed decisions in the design process.